OpenStudy (anonymous):

What is the solution to the equation = 1 over the square root of 8= 4^(m – 2) ?

Parth (parthkohli):

\[\log_6 x + \log_6(x + 1) = 1 \]Now use the property:\[ \log_na + \log_nb = \log_nab\]

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i messed up by accident ha

OpenStudy (anonymous):

\[\frac{1}{\sqrt{8}}=2^{-\frac{3}{2}}\]

Parth (parthkohli):

\[\left((2^3)^{1 \over 2}\right)^{-1}= (2^2)^{m - 2} \]

Parth (parthkohli):

\[2^{-1.5} = 2^{2m - 4} \]Now you can equate the exponents.

OpenStudy (anonymous):

swaggg , thanks .

Parth (parthkohli):

You're welcome.

Parth (parthkohli):

Dirty answer, isn't it?